A firm has the following demand function 𝑷 = πŸ”πŸŽ βˆ’ 𝟎. πŸ“π‘Έ and its total cost are


defined by 𝑻π‘ͺ = πŸπŸ‘ + 𝑸.




a. Find the maximum revenue.



b. Find the production to optimise the profit.



c. Verify if that the marginal revenue and marginal cost are the same at


the profit maximising production level.

1
Expert's answer
2022-04-07T13:07:29-0400


a) Maximum revenue is determined from the equation TR'=0. TR=PQ=(60-0.5Q)Q=60Q-0.5Q2

TR'=60-Q=0; Q=60

b) Optimal profit is determined from the equation MR=MC. MR=TR'=60-Q; MC=TC'=(13+Q)'=1; 60-Q=1; Q=59

c) At Q=59 MR=60-59=1=MC